OpenStudy (anonymous):

Solve: 6(x - 2) + 3(x - 2) = 18. i need some tutor help

mathslover (mathslover):

\[\huge{6x-12+3x-6=18}\] \[\huge{9x-15=18}\] \[\huge{9x=23}\] \[\huge{x=\frac{23}{9}}\]

OpenStudy (anonymous):

need help

OpenStudy (anonymous):

no is

OpenStudy (anonymous):

i think its 22 over 9

mathslover (mathslover):

\[\huge{x=\frac{9}{2}}\]

OpenStudy (anonymous):

my answer choices are A. x = 22/9 B. x = 2 C. x = 4

mathslover (mathslover):

c. x=4

mathslover (mathslover):

\[\huge{6(x - 2) + 3(x - 2) = 18}\] \[\huge{6x-12+3x-6=18}\] \[\huge{9x=36}\]\[\huge{x=4}\]

mathslover (mathslover):

have to go ..